OpenStudy (anonymous):

Find the roots of the polynomial equation: x^2 – 8x – 48 = 0

OpenStudy (anonymous):

answer a) -6,8 b) -4, 12 c)-12, 4 d)-8,6

OpenStudy (anonymous):

b

OpenStudy (anonymous):

x^2-8x-48 = (x-12)(x+4) = 0; Therefore, when x = 12 or x=-4 the equation is true.